OpenStudy (anonymous):

2/3!+4/5!+6/7!+....to infinity is equal to?

OpenStudy (anonymous):

What you've got there is a basic series...\[\sum_{n=0}^{\infty}\frac{n+2}{n!}\]Maybe you should try using the ratio test. That's what I'd first use to see if it converges.

OpenStudy (anonymous):

the series notation is wrong

OpenStudy (anonymous):

i dont think so buddy

OpenStudy (anonymous):

Sorry, the sum's off. :P Don't have time to work it out right now, but I'm interested to see if anyone's got it to a concise form

OpenStudy (anonymous):

m with you..;-)

OpenStudy (anonymous):

1/2 is the answer?

OpenStudy (anonymous):

no its 1/e

OpenStudy (anonymous):

which book r u using?

OpenStudy (anonymous):

"rudiments of mathematics" its a book for wbjee(west bengal joint entnc xm)

OpenStudy (anonymous):

Wow, I wasn't even thinking. Try this\[\sum_{n=1}^{\infty}{\frac{2n}{(2n+1)!}}\]

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

(2n+2)/(2(n+3) *(2n+1)!/2n

OpenStudy (anonymous):

(2n+2)/(2n+3)! *(2n+1)!/2n

OpenStudy (anonymous):

wow so u got the expression right finally

OpenStudy (anonymous):

\[2n+2 \over {(2n+2)(2n+3)(2n)}\]

OpenStudy (anonymous):

\[\lim n \rightarrow \infty [1/2n(2n+3)\]

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

This just tell us that it converge but not to what

OpenStudy (anonymous):

yup,convergent

OpenStudy (anonymous):

1/e= e^-1[1+1/2! +1/3!...]^-1 that can be expanded using binomial expension

OpenStudy (anonymous):

\[[1+(1/2/!+1/3!...)]^{-1}=1- (1/2!+1/3!...)+(1/2!+ 1/3!...)^2....\]

OpenStudy (anonymous):

how come the sum is 2/3!+4/5!+6/7!+....?

OpenStudy (anonymous):

i dont kno. thats what the book says.. its wrong isn't it?

OpenStudy (anonymous):

seems to be, not sure

OpenStudy (anonymous):

wow..great thinkin there buddy.. thanx!